# Robotics meets Fluid Dynamics: A Characterization of the Induced Airflow around a Quadrotor

### v1

@@ IV-B Quadrotors
For experiments three different quadrotors are used whose properties are summarized in Tab. I. The vehicles Offboard 1 and Offboard 2 are research drones that can autonomously fly missions [47] and are controlled through the motion-capture system. Both drones share the same frame, however drone 2 has double the mass compared to drone 1. We also use a DJI Matrice 300 drone as a much larger and heavier vehicle. This allows us to demonstrate the generality of the proposed jet flow model and underlying theoretic framework. In contrast to the research drones, the commercial DJI drone in this study must be flown manually throughout the entire experiment.
@@
Mass
Propeller Diameter
Motor Distance
Induced Velocity
Offboard 1
0.572 kg
12.95 cm
26.6 cm
6.66 m/s
Offboard 2
1.207 kg
12.95 cm
26.6 cm
9.66 m/s
Matrice 300
6.300 kg
53.34 cm
89.4 cm
5.36 m/s
IV-C Data Collection
To collect the experimental data, each vehicle is flown in a grid-like pattern where the drone approaches a point, steadily hovers for 5 s to allow the flow to (re-)develop and then slowly translates to the next point in the three-dimensional flight path. For the autonomous drones we sample a coarse grid spanning $3\times 3\times 2.5\,\mathrm{m}$ with a spacing of 15 cm in the $xy$-plane and 50 cm in $z$ and a fine grid spanning $0.5\times 0.5\times 2.5\,\mathrm{m}$ with a resolution of 10 cm in the $xy$-plane and 30 cm in $z$. For the Matrice 300 the pilot covered a much larger area of $5\times 6\times 4.5\,\mathrm{m}$.
Transforming the flight data into the drones’ reference frame, we filter out all data points where the vehicle moves with more than 0.1 m/s speed during the sampling interval. This reduces the influence of transient effects and ensures high data quality. Despite being in a large indoor space, the anemometer registers a small ambient flow in the range of 0.06 m/s to 0.12 m/s. As this is flow is already observed before the drone takes off, it is not primarily caused by recirculation effects but related to the testing facility. The background level is about 2 orders of magnitudes below the induced velocities. We correct for the static offset by subtracting the ambient airflow speed from the measurements.
After filtering we obtain 12000 s of flight data for each of the three vehicles. The raw measurement data is in the form of a position measurement and a corresponding anemometer measurement. For further processing, this scattered data is converted into binned gridded data by combining all measurements within a grid cell. To increase the robustness of the measurements, the median value of the cell is used and the empirical standard deviation is calculated.
@@ V Results: Near Field
Figure 4: Visualization of the near field of (a) the Offboard 1 drone and (b) the Matrice 300. The length and velocity scales are normalized. The influence of the individual propellers is clearly visible at $z=0$ and diminishes at a normalized distance of about 2.
Figure 5: The plot shows the distance $\tilde{\delta}$ of the flow tube centerline to the $\mathbf{z}_{\mathcal{B}}$ axis. Very close to the rotor-plane of the vehicle (left side of the plot) the flow has not fully developed. From between two and three length scales below the drone the flow has fully developed with the highest velocity being measured on the negative z-axis of the drone.
Figure 6: Radial velocity profiles for the three different drones. The colors show different height levels. The shaded area indicates the $1\sigma$ uncertainty region around the mean, the line shows the median value. The dashed line shows the induced velocity at hover.
In a first analysis we focus on the flow directly below the drone in the region where the flow contributions of the four individual propellers are seperated and not yet have fully merged into one turbulent jet. The goal is to analyse how the flow from the individual propellers merge into a combined turbulent jet for the different quadrotors.
Figure 4 visualizes the flow field for a $yz_{\mathcal{B}}$-slice from the measurement data. The quadrotor is located at the ‘top’ of the plot (e.g. at $\tilde{z}$=0) and one can clearly see the two separate jets induced by the left and the right propellers. The Offboard 1 and Matrice 300 drone show similar flow patterns as we normalize the axes with the drones’ tip-to-tip dimensions. This indicates that the tip-to-tip length is an appropriate scaling for the drones’ flow geometry.
In between the two jets, a region of reduced flow speed can be observed. This depression is due to the separation of the drones’ rotors and the blocking of the flow by its main body. Such flow structures have previously been observed in CFD simulations [14] and validation studies [13], and are consistent with our results. Afterward, the jets start to converge at about one length scale below the rotor plane. From about two length scales below the drone, the jets appear to have fully merged. In metric coordinates this corresponds to about 0.5 m below the Offboard 1 drone and 2 m below the Matrice 300.
Figure 5 quantitatively describes the merging of the individual jets. The distance of the maximum propeller flow velocity centerline to the drones’ negative $\mathbf{z}_{\mathcal{B}}$ axis is considered. This distance is estimated by first extracting the $\tilde{y}$-axis velocity profile for different distances below the drone. Then this empirical profile is approximated with a $z_{\mathcal{B}}$-axis symmetric bimodal Gaussian where the means are shifted ±$\pm\tilde{\delta}$ along the $\tilde{y}$-axis. In agreement with related works [14] we find that for all drones in this study, the individual rotor flows merge between 2 and 2.5 tip-to-tip dimensions below the drone.
@@ VI Results: Far Field
From the preliminary considerations of the airflow close to the vehicle, we see that the individual contributions from the quadrotor’s four propellers merge together about two length scales below the drone. In this section, we present our novel experimental findings to answer the question:‘How much wind do the quadrotors make away from their propellers in the far field?’.
@@
Figure 7: Radial velocity profiles for the three drones at a normalized height of $\tilde{s}=-3.0$. The shaded area indicates the $1\sigma$ uncertainty region around the mean, the line shows the median value. The proposed normalization leads to all curves being similar within measurement accuracy. The black line represents the radial velocity profile for a turbulent jet eq. (2).
Figure 8: Downstream scaling parameters of the flow. Only datapoints in the far field region (marked with a circle) are used to fit the function predicted by turbulet jet theory. (a) The plot depicts the (normalized) tube half-width $\tilde{r}_{1/2}$ as a function of the (normalized) distance to the rotor plane $\tilde{s}$. We recover the characteristic spreading angle for turbulent jets of 12 deg. (b) The plot focuses on the (normalized) centerline velocity $\tilde{U}_{\text{C}}$ as a function of the (normalized) distance to the rotor plane. We recover the characteristic $1/(s-s_{0})$ scaling from eq. (4).
Figure 9: Unified formulation of the flow. The measurement samples from all drones across the entire far field show good agreement with turbulent jet theory eq. (2) (black line).
@@ VI-A Radial Velocity Profiles
Figure 6 shows radial profiles of the mean velocity for the three drones at different heights. First, for all drones, the peak measured velocities exceed the calculated induced velocity at hover by 25 % at maximum. This small mismatch with the calculated propeller induced flow is expected as the velocity profile in the merged flow is not constant across the cross section and by conservation of momentum the peak velocity is higher. Second, we compare measurements at different $s$-coordinates for each drone and observe that the mean flow velocities all decay radially from the drone. Third, the downstream peak velocity rapidly decreases away from the drone, while the flow domain spreads out. Finally, comparing measurements above and below the drone, we see that the flow velocity above the drone ($s=-0.35\,\mathrm{m}$ and $s=-0.65\,\mathrm{m}$) is much lower compared to below the drone as the flow is uniformly sucked into the propellers from all directions.
We now look at the same data using the normalization introduced in Sec. III-D. The plot in Fig. 7 corresponds to a normalized $\tilde{s}$ level at 3 length scales (e.g. tip-to-tip distance) below the drone. The normalized, time-averaged measured velocity $\langle\tilde{U}\rangle$ is plotted as a function of the normalized radial coordinate $\tilde{r}$ and each color now corresponds to one drone. We can clearly see that, within the given measurement accuracy, the curves overlap which empirically validates the proposed normalization as it makes the airflow around different vehicles similar. Additionally, the radial velocity profile across all drones is well captured fitting eq.(2), inset in black in Fig. 7. This provides the first evidence that the combined rotor flow can be treated as a turbulent jet.
@@ VI-B Jet Flow Scaling
Next, we focus on the expansion of a turbulent jet which is described by eq. (5). For a turbulent jet, the flow expands linearly as a cone with approximately 12 deg opening angle [24] over a wide range of flow Reynolds number. Furthermore, the jet’s centerline velocity scales inversely proportional with the distance to the nozzle as described by eq. (4).
In Figure 8 we find very good agreement with these two characteristics for the combined rotor flow of a quadrotor after $\tilde{s}=2$. Figure 8a and b show the normalized half-width $\tilde{r}_{1/2}$ and the normalized centerline velocity $\langle\tilde{U}_{\text{C}}\rangle$ of the radial velocity profiles as a function of the normalized distance $\tilde{s}$ to the drone’s rotor plane. The half-width and velocity have been obtained by fitting the velocity profile eq. (2) to the measurements at each $\tilde{s}$-slice for each drone. The triangles indicate datapoints located in the near-field (not considered for the fit) and circles indicate measurements from the far field (considered for the fit). The black lines represent a fit using the function form prescribed by the turbulent jet model eq. (5) and (4). The parameters obtained from the fit are
$B\,d=6.84\,,\quad S=0.0997\,\text{~{}and}\quad s_{0}=3.08\;.$ (8)
The scaling parameters (8) can be used to fully describe the turbulent jet flow in eq. (2). This velocity profile from eq. (2) is shown with the black line in Fig. 9 and is not fitted to the data. The measurement samples are taken from all drones across the entire far field and measurement scaling is computed using eq. (5) and (4) with the parameters above. The good agreement with the theory shows that the measured velocity of a quadrotor field in both axial and radial dimension is well captured by the theory of turbulent jets.
@@ VI-C Unified Model
    Normalization parameter: Use the physical parameters (mass, propeller size, number of propellers) of the quadrotor to calculate the induced velocity at hover $U_{\rm{H}}$ using eq. (1) and the tip-to-tip length $l$
@@
VII Conclusion and Discussion
This work presents a model for the far field flow of a drone, based on classical turbulent jet theory. We have recorded a large-scale dataset comprising three drones and containing over 9 h of flight data in a motion-capture system while measuring the airflow with a hot-ball flow probe. Leveraging normalization techniques relying on the physical parameters of the vehicle (mass, propeller size, vehicle size), the airflow of different sizes of drones is comparable.
@@ VII Conclusion and Discussion
We study the flow in the near-field flow close to the drone and find that the individual flows from the propellers merge approximately two tip-to-tip length scales below the drone. In this near-flow region, the interactions between the flow and the drone are dominant which can only be fully captured through CFD simulations. However, when considering the effects of the induced flow on the environment and other agents, the primary concern is the far-field extending from about two length scales downwards.
@@
Through the use of a scaling approach from turbulent jet theory, we develop a unified model that describes the far field flow below the drone. The experiments show that, despite the simplicity of the model, it is able to accurately describe the flow. This makes our model ideally suited for integration in a dynamic planner and in multi-agent scenarios. We believe that this is an important step towards safer and less intrusive drones, an aspect becoming increasingly important with the increasing adoption of quadrotors. Understanding the flow also enables optimized sensor placement for scientific applications and potentially higher efficiency when quadrotors are deployed in an agricultural context.

### v2

@@ IV-B Quadrotors
For experiments, six different quadrotors are used whose properties are summarized in Tab. I. Kolibri, Offboard 1 and Offboard 2 are research drones that can autonomously fly missions [46] in the motion-capture system. Both Offboard drones share the same frame, however, drone 2 has double the mass compared to drone 1. To demonstrate the generality of the proposed jet flow model we also use a DJI Matrice 300 drone as a much larger and heavier vehicle. The aforementioned drones have an uncanted propeller arrangement, meaning that their axis of rotation is parallel to the $z_{\mathcal{B}}$-axis. We additionally use two commercial drones, respectively a DJI Mavic 3E (enterprise) and a Flyability Elios 3, with canted propeller configurations to better understand the limitations of our model. The commercial DJI and Flyability drones must be flown manually throughout the entire experiment.
IV-C Data Collection
To collect the experimental data, each vehicle is flown in a grid-like pattern where the drone approaches a point, steadily hovers for 5 s to allow the flow to (re-)develop, and then slowly translates to the next point in the three-dimensional flight path. For the autonomous drones, we sample a coarse xy-grid spanning ${10\times 10}$ length-scales (resolution 0.66 $l$) and a fine xy-grid spanning ${3\times 3}$ length-scales (resolution 0.33 $l$). In $z$ we cover a range of up to 3m above the probe. For the Mavic 3E and the Elios 3 the pilot sampled a similar pattern to the autonomous drone. For the Matrice 300 the pilot covered an area amounting to $5\times 6\times 4.5\,\mathrm{m}$.
We filter out all data points where the vehicle moves with a speed greater than 0.1 m/s to reduce the influence of transient effects. Despite being in a large indoor space, the anemometer registers a small ambient flow in the range of 0.06 m/s to 0.12 m/s. As this flow is already observed before the drone takes off, it is not primarily caused by recirculation effects but related to the testing facility. The background level is about 2 orders of magnitudes below the induced velocities and we correct for this static offset by subtracting the ambient airflow speed from the measurements.
After filtering we obtain around $10000\,\mathrm{s}$ of flight data (position and corresponding anemometer measurement) for each of the vehicles. For further processing, this scattered data is binned into gridded data as the median of all measurements within a grid cell.
@@ IV-C Data Collection
Mass Propeller Cant Propeller Diameter Motor Distance Induced Velocity
Kolibri 0.230 kg uncanted 7.37 cm 11.8 cm 7.41 m/s
Offboard 1 0.572 kg uncanted 12.95 cm 26.6 cm 6.66 m/s
Offboard 2 1.207 kg uncanted 12.95 cm 26.6 cm 9.66 m/s
Matrice 300 6.300 kg uncanted 53.34 cm 89.4 cm 5.36 m/s
Mavic 3E 0.958 kg inward 23.88 cm 38.5 cm 4.67 m/s
Elios 3 2.398 kg outward 12.70 cm 27.5 cm 13.89 m/s
@@ V Results: Near Field
In a first analysis, we focus on the flow directly below the drone. In this region the flow contributions of the four individual propellers are separated and have not yet fully merged into a combined drone downwash. The goal is to analyse how the individual propeller flows eventually merge into one turbulent jet for the different quadrotors.
Figure 4 exemplarily visualizes the flow field for a $yz_{\mathcal{B}}$-slice from the measurement data. The quadrotor is located at the ‘top’ of the plot (e.g. at $\tilde{z}$=0) and one can clearly see the two separate flows induced by the left and the right propellers. The Offboard 1 and Matrice 300 drone show similar flow patterns as we normalize the axes with the drones’ motor-distance $l$. This indicates that the motor-distance is an appropriate scaling for the drones’ flow geometry.
In between the two jets, a region of reduced flow speed can be observed. This depression is due to the separation of the drones’ rotors and the blocking of the flow by its main body. Such flow structures have previously been observed in CFD simulations [13] and validation studies [12], and are consistent with our results. Afterwards, the jets start to converge at about one length scale below the rotor plane. From about 2.5 length scales below the drone, the jets appear to have fully merged.
Figure 4: Visualization of the near field of (a) the Offboard 1 drone and (b) the Matrice 300. The length and velocity scales are normalized. The influence of the individual propellers is clearly visible at $z=0$ and diminishes at a normalized distance of about 2.5 as indicated by the dotted white line.
Figure 5: Distance $\tilde{\delta}$ of the flow tube centerline to the $\mathbf{z}_{\mathcal{B}}$ axis. Close to the rotor-plane of the vehicle (left side) the flow has not fully developed. From 2 to 3 length scales below the drone the flow fully develops with the highest velocity being measured on the negative z-axis of the drone.
Figure 5 quantitatively describes the merging of the individual jets. The distance of the maximum propeller flow velocity centerline to the drones’ negative $\mathbf{z}_{\mathcal{B}}$ axis is considered. This distance is estimated by first extracting the $\tilde{y}$-axis velocity profile for different distances below the drone. Then this empirical profile is approximated with a $z_{\mathcal{B}}$-axis symmetric bimodal Gaussian where the means are shifted ±$\pm\tilde{\delta}$ along the $\tilde{y}$-axis. In agreement with related works [13] we find that for our drones the individual rotor flows merge between 1.5 and 2.5 motor-distances below the drone.
@@ VI Results: Far Field
From the preliminary considerations of the airflow close to the vehicle, we see that the individual contributions from the quadrotor’s four propellers merge about 2.5-length scales below the drone. In this section, we present our novel experimental findings to answer the question:‘How much wind do the quadrotors make away from their propellers in the far field?’.
@@
Figure 6: Radial velocity profiles for the Kolibri drone. The colors represent different height levels. The shaded area indicates the $1\sigma$ uncertainty region around the mean, the line shows the median value.
Figure 7: Radial velocity profiles for the drones with a planar propeller layout at a normalized height of $\tilde{s}=3.0$. The shaded area indicates $1\sigma$ uncertainty around the mean, the line shows the median value. The proposed normalization leads to all curves being similar in measurement accuracy. The black line represents the radial velocity profile for a turbulent jet eq. (3).
Figure 8: Downstream scaling parameters of the flow. Only data points in the far field region (marked with a circle) are used to fit the function predicted by turbulent jet theory. (a) The plot depicts the (normalized) tube half-width $\tilde{r}_{1/2}$ as a function of the (normalized) distance to the rotor plane $\tilde{s}$. We recover a characteristic spreading angle of 10 deg. (b) The plot focuses on the (normalized) centerline velocity $\tilde{U}_{\text{C}}$ as a function of the (normalized) distance to the rotor plane. We recover the characteristic inverse $Bd/(s-s_{0})$ scaling from eq. (5). The parameters for (5) and (6) are fitted jointly to the data. (c) A visualization of the measured Kolibri drone flow field. The tube half-width from (a) is overlaid in white.
@@ VI-A Radial Velocity Profiles
Figure 6 exemplarily shows radial profiles of the mean velocity for the Kolibri drone at different heights. First, the measured peak velocities exceed the calculated induced velocity at hover by 25 % at maximum. This mismatch with the calculated propeller-induced flow is expected as the velocity profile in the merged flow is not constant across the cross-section and by conservation of momentum the peak velocity is higher. Second, we compare measurements at different $s$-coordinates and observe that the mean flow velocities all decay radially from the drone. Third, the downstream peak velocity rapidly decreases away from the drone, while the flow domain spreads out.
We now look at the data for multiple drones using the normalization introduced in Sec. III-D. The plot in Fig. 7 corresponds to a normalized $\tilde{s}$ level at 3 length scales (motor-distance) below the drone. The normalized, time-averaged measured velocity $\langle\tilde{U}\rangle$ is plotted as a function of the normalized radial coordinate $\tilde{r}$ and each color now corresponds to one drone. We can clearly see that, within the given measurement accuracy, the curves overlap which empirically validates the proposed normalization as it makes the airflow around different vehicles similar. Additionally, the radial velocity profile across all drones is well captured fitting eq.(3), inset in black in Fig. 7. This provides evidence that the combined rotor flow can be treated as a turbulent jet.
@@ VI-B Jet Flow Scaling
Next, we focus on the expansion of a turbulent jet which is described by eq. (6). For a turbulent jet, the flow expands linearly as a cone with approximately 12 deg opening angle [23]. Furthermore, the jet’s centerline velocity scales inversely proportional to the distance to the nozzle as described by eq. (5).
In Figure 8 (a) and (b) we find good agreement with these two key characteristics for the combined rotor flow of a quadrotor after $\tilde{s}=2.5$. Figure 8a and b show the normalized half-width $\tilde{r}_{1/2}$ and the normalized centerline velocity $\langle\tilde{U}_{\text{C}}\rangle$ of the radial velocity profiles as a function of the normalized distance $\tilde{s}$ to the drone’s rotor plane. The half-width and velocity have been obtained by fitting the velocity profile eq. (3) to the measurements at each $\tilde{s}$-slice for each drone. The triangles indicate datapoints located in the near-field (not considered for the fit) and circles indicate measurements from the far field (considered for the fit). The black lines represent a fit using the function form prescribed by the turbulent jet model eq. (5) and (6). The parameters obtained from the fit are
$B\,d=10.11\,,\quad S=0.07668\,\text{~{}and}\quad s_{0}=-5.817$ (9)
for the unified flow. Note that only drones with uncanted propeller configurations are used to obtain the fit. Figure 8c shows visualization of the Kolibri far-field velocity profile with the white lines indicating the half-width (see Fig. 8a).
The scaling parameters (9) can be used to fully describe the turbulent jet flow in eq. (3). This velocity profile from eq. (3) is shown with the black line in Fig. 9 and is not fitted to the data. The measurement samples are taken across the entire far field and measurement scaling is computed using eq. (5) and (6) with the parameters above. The good agreement with the theory shows that the measured velocity of a quadrotor field in both axial and radial dimension is well captured by the theory of turbulent jets.
Figure 9: Unified formulation of the flow. The measurement samples from all planar drones across the entire far field show good agreement with turbulent jet theory eq. (3) (black line). Drones with canted propellers show a slightly narrower or a wider jet, depending on the geometry.
We note that the measured velocities in Fig. 9 for uncanted drone designs are well-centered around the theoretical curve in the region for low $\xi$ but seem to fall below the theory curve $U_{\text{theory}}$ eq. (3) around ${\xi>2}$ until the measurement noise dominates from ${\xi>5}$ onwards. To corroborate this visual impression, a one-sided $t$-test for the mean of the distribution of measurement residuals ${\epsilon=U_{\text{theory}}-\langle U\rangle/U_{C}}$ is conducted. We verified the normality of the residuals with a $\chi^{2}$ goodness-of-fit test. The $t$-test results (see Tab. II) show that for ${\xi>2}$ the theory overpredicts the velocity. This is a well-known result for turbulent jets [23]. However, in practice it can be safely ignored due to the negligible magnitude of the mismatch w.r.t typical ambient airflow.
TABLE II: Hypothesis Test ($\alpha=0.05$):   $H_{0}:\mu(\epsilon)=0$,   $H_{A}:\mu(\epsilon)>0$
Interval               $\xi\in$ [0  1) [1 2) [2  3) [3  4) [4  5) [5  6)
Reject $H_{0}$, e.g., $\mu(\epsilon)>0$ False False True True True False
In Fig. 9 the two drones with canted propeller planes are also included in the plot and represented with squares and triangles. The Mavic 3E design features an inward cant, effectively focusing the flow. This is consistent with the observation that the jet is narrower with slightly higher centerline velocities. The Elios 3 on the other hand has outwardly canted propellers which leads to a wider jet with lower centerline velocities but a larger half-width. This effect is especially notable around ${\xi=0}$ and ${\xi=2.5}$.
@@ VI-C Unified Model
    Normalization parameter: Use the physical parameters (mass, propeller size, number of propellers) of the quadrotor to calculate the induced velocity at hover $U_{\rm{H}}$ using eq. (1) and the motor distance $l$.
@@
VII Controller Integration
To highlight that our turbulent-jet model is relevant to real-world multi-agent scenarios, we conclude our work by demonstrating how the model can be used to compute a feed-forward downwash compensation in a scenario where one drone needs to pass below another drone.
VII-A Downwash Compensation
When flying below another drone, the lower drone needs to additional power to maintain hover thrust. Similar to eq. (1), momentum theory can be used to calculate the airspeed $U_{H}^{\prime}$ below the propeller. However, in this case the air above the propeller is not at rest but moving downwards with $U_{D}$ [28],
$U_{H}^{\prime}=\frac{U_{D}}{2}+\sqrt{\left(\frac{U_{D}}{2}\right)^{2}+U^{2}_{H}}\;,$ (10)
where $U_{H}$ from is the induced velocity at hover in still air from eq. (1), and $U_{D}$ is calculated according to the unified model presented in the previous section. Both $U_{D}$ and $U^{\prime}_{H}$ depend on the relative position $\Delta\mathbf{p}$ of the drones. The total aerodynamic power required to hover inside the downwash is ${P_{H}^{\prime}=U_{H}^{\prime}\cdot T_{H}}$. Introducing the relative induced velocity ${\alpha=U_{D}/U_{H}}$ we get the relative power $\beta$ as
$\beta:=\frac{P_{H}^{\prime}}{P_{H}}=\frac{\alpha}{2}+\frac{\sqrt{\alpha^{2}+4}}{2}\;.$ (11)
The electronic speed controllers (ESC) on the Kolibri drone exhibit a quadratic relationship between the throttle command and the power consumption. Consequently, the throttle setpoint computed by the controller is scaled with $\sqrt{\beta}$ before sending it to the vehicle’s motor controllers. This scaling factor is time-varying as it depends on $\Delta\mathbf{p}$. We implement eq. (11) on top of an MPC controller [46].
VII-B Experimental Results
To demonstrate the effectiveness of the control scheme, we perform the following experiment: the Offboard 2 drones hovers at 3 m height and the much smaller Kolibri passes 1 and 2 meters exactly below at a translation speed of 1 length-scale per second. The results are shown in Fig. 10 and Table III and one can see that the baseline controller without the proposed downwash correction has a much higher altitude tracking error compared to the corrected controller. Comparing the RMSE in height we see a 3 to 5-fold improvement with the proposed controller and observe that the performance improvement is larger in stronger downwash (e.g., closer to the upper drones).
Figure 10: Experiment showing that our downwash model can be used to improve the control performance in a multi-drone scenario. The dotted line starting at 2.5 length scales below Offboard 2 indicates the half-width calculated from eq. (6). The depicted drone is true-to-scale in width. The right side shows a time-lapse photograph of the run at 2 m.
TABLE III: Height Tracking Performance of the Controller
Controller Height [m] RMSE [mm] Mean Err. [mm]
Without Correction 1 m 69.1 -37.6
With Correction 1 m 24.8 18.9
Without Correction 2 m 134.2 -87.1
With Correction 2 m 25.4 8.1
VIII Conclusion and Discussion
@@
This work presented a model for the far field flow of a drone, based on classical turbulent jet theory. We recorded a large-scale dataset comprising six drones and containing over 16 h of flight data in a motion-capture system while measuring the airflow with a hot-ball flow probe. We studied the flow in the near-field flow close to the drone and found that the individual flows from the propellers merge approximately 2.5 motor-distance length scales below the drone. In this near-flow region, the interactions between the flow and the drone are dominant which can only be fully captured through CFD simulations. However, when considering the effects of the induced flow on the environment and other agents, the primary concern is the far-field extending from about two length scales downwards.
Through the use of appropriate scaling laws for turbulent flows, we developed a unified model that describes the far field flow below the drone as a turbulent jet. The experiments show that, despite the simplicity of the model, it accurately describes the flow for a wide range of drones at hover. Jet flows are well-established in fluid mechanics. For example, in the case of cross-flows [47], jets include empirical scalings for the center-line shape, relevant to vertical and forward flight [28], and rotor induced-vortex flows [48]. Moreover, experiments on impinging jets [49] and similarity-solutions for wall-bounded jet flows [50] help understand modulations in ground effect and near boundaries as a first step towards confined environments with internal circulation. Showing that the induced flow of a drone is a turbulent jet, we believe that applying the above results from fluid mechanics presents an interesting avenue for further research.
@@
Finally, we demonstrated that the accuracy and efficiency make our model ideally suited for integration in a controller for multi-agent scenarios. We believe that this is an important step towards safer and less intrusive drones, an aspect becoming increasingly important with the increasing adoption of quadrotors. Understanding the flow also enables optimized sensor placement for scientific applications and potentially higher efficiency when quadrotors are deployed in an agricultural context.